# Pan-LUT: Efficient Pan-sharpening via Learnable Look-Up Tables

### v1

@@ 3.2 Spectral Transformation
After locating 32 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$O_{(x,y,z,s,e)}=PInterpolation(LUT[L],{\{\mathbf{d}_{l}\}}),$ (4)
where $PInterpolation(\cdot)$ denotes the pentalinear interpolation.
3.3 Spatial Details Transformation
PGLUT is essentially a channel-wise 5D LUT that operates globally, which limits its ability to capture local spatial information. To effectively capture fine-grained spatial details and adaptively learn local contexts, we propose the Spatial Details Lookup Table (SDLUT) and the Adaptive Aggregation Lookup Table (AALUT). As illustrated in Figure 2, given a pixel $p_{(w,h)}$, SDLUT processes this pixel along with its neighboring pixels as input. During the training phase, we employ a rotation ensemble strategy to further expand the receptive field, which can be formulated as:
$\begin{gathered}{V_{1}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w,h+1)},p_{(w+1,h+1)}),\\ {V_{2}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w+1,h-1)},p_{(w,h-1)}),\\ {V_{3}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w,h+1)},p_{(w+1,h+1)}),\\ {V_{4}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w,h+1)},p_{(w+1,h+1)}).\\ \end{gathered}$ (5)
Subsequently, a learnable AALUT is introduced to adaptively integrates the four outputs. Specially, the AALUT takes the output from SDLUT as input and combines the spatial information to enhance the spatial representation:
${V}_{as}=f_{AALUT}({V_{1}},{V_{2}},{V_{3}},{V_{4}}),$ (6)
where $f_{AVGLUT}(\cdot)$ denotes the lookup and interpolation process in the LUT retrieval.
@@ 3.4 Loss Function
To achieve satisfying pan-sharpening results, we propose a joint loss for network training. We first utilize the L1 loss:
$\mathcal{L}_{1}=\left\|HRMS-GT\right\|_{1},$ (7)
@@
$\mathcal{L}_{s}=\mathcal{L}_{s}^{PG}+\mathcal{L}_{s}^{SD}+\mathcal{L}_{s}^{AA},$ (8)
$\mathcal{L}_{m}=\mathcal{L}_{m}^{PG}+\mathcal{L}_{m}^{SD}+\mathcal{L}_{m}^{AA},$ (9)
where $\mathcal{L}_{s}^{PG}$, $\mathcal{L}_{s}^{SD}$, and $\mathcal{L}_{s}^{AA}$ denote the smoothness regularizations for PGLUT, SDLUT, and AALUT, while $\mathcal{L}_{m}^{PG}$, $\mathcal{L}_{m}^{SD}$, and $\mathcal{L}_{m}^{AA}$ represent the monotonicity regularizations for PGLUT, SDLUT, and AALUT, respectively.
Taking AALUT as an example, the smoothness regularization can be defined as:
$\begin{gathered}\mathcal{L}_{s}^{AA}=\sum_{O\in\{l,o,c,a\}}\sum_{i,j,k,m=0}^{N-1}(\left\|O_{(i+1,j,k,m)}-O_{(i,j,k,m)}\right\|^{2}\\ +\left\|O_{(i,j+1,k,m)}-O_{(i,j,k,m)}\right\|^{2}\\ +\left\|O_{(i,j,k+1,m)}-O_{(i,j,k,m)}\right\|^{2}\\ +\left\|O_{(i,j,k,m+1)}-O_{(i,j,k,m)}\right\|^{2}),~{}\end{gathered}$ (10)
where $N$ represents the number of bins in each dimension of the LUT. $O_{(i,j,k,m)}$ is the corresponding output for the defined sampling point $(i,j,k,m)$ in LUT. The definitions of $\mathcal{L}{s}^{PG}$ and $\mathcal{L}{s}^{SD}$ are similar to those in Eq. 10.
The monotonicity regularization in AALUT can be defined as:
$\begin{gathered}\mathcal{L}_{m}^{AA}=\sum_{O\in\{l,o,c,a\}}\sum_{i,j,k,m=0}^{N-1}[g(O_{(i,j,k,m)}-O_{(i+1,j,k,m)})\\ +g(O_{(i,j,k,m)}-O_{(i,j+1,k,m)})\\ +g(O_{(i,j,k,m)}-O_{(i,j,k+1,m)})\\ +g(O_{(i,j,k,m)}-O_{(i,j,k,m+1)})],~{}\end{gathered}$ (11)
where $g(\cdot)$ denotes the ReLU activation function. Similarly, $\mathcal{L}_{m}^{PG}$ and $\mathcal{L}_{m}^{SD}$ are defined in the same way as in Eq. 11.
@@ 4 Experiments
In this section, we provide a comprehensive overview of the experimental setup, including the evaluation datasets, performance metrics, and implementation details. We then present both quantitative and qualitative comparisons with other methods. Additionally, we conduct ablation studies to validate the contribution of each component.
@@ 4.1 Datasets
Remote sensing datasets from three satellites are used in our experiments, including WorldView-II (WV2), GaoFen2 (GF2) and WorldView-III (WV3). Due to the absence of high-resolution multispectral ground truth images in these datasets, we generate the training set using the Wald protocol tool [36]. Specifically, given the original MS image and its corresponding high-resolution PAN image, they are downsampled by a factor of $r$ to obtain image pairs of MS and PAN, with $r$ set to 4. During training, the original high-resolution MS image is treated as the ground truth, while the MS and PAN images serve as the input image pairs.
@@
$D_{\lambda}\downarrow$ 0.1026 0.1110 0.1057 0.1063 0.0987 0.1034 0.1042 0.0966 0.0967 0.0966  0.0953 0.0962
$D_{S}\downarrow$ 0.1409 0.1556 0.1446 0.1443 0.1312 0.1305 0.1476 0.1274 0.1271 0.1272  0.1129 0.1253
@@
4.2 Implementation Details
We compare the proposed Pan-LUT against several pan-sharpening methods on reduced-resolution scenes from WV2, WV3, and GF2 datasets. Specifically, we choose four traditional pan-sharpening techniques: Brovey [11], IHS [2], SFIM [29] and GS [21], along with ten deep learning-based approaches: PNN [32], MSDCNN [50], Pan-GAN [31], SFDI [60], UCGAN [58], PanFlow [48], PSCINN [37], Pan-Mamba [13] and TA-DiffHQP [41].
@@ 4.2 Implementation Details
4.3 Comparison with Other Methods
Evaluation on Reduced-resolution Scene. The quantitative results across three datasets are presented in Table 1, with the best results highlighted in red. Compared to traditional methods, our approach achieves an average PSNR improvement of 4dB, 5dB, and 6dB across the three datasets, while maintaining inference speeds comparable to those of traditional methods. Compared to DNN-based methods, our method exhibits an average PSNR performance decrease of 2 dB, 4 dB, and 2 dB across the three datasets. However, as shown in the table, only PNN is capable of processing 4K-resolution images. Notably, our method is 181.97 times faster than PNN, while the performance gap remains minimal at 0.92 dB, 0.46 dB, and 1.12 dB across the three datasets. It is worth noting that the proposed Pan-LUT does not incorporate any network structure, yet it outperforms some DNN-based methods, such as Pan-GAN and UCGAN, in terms of both performance and inference time. We also provide visual comparisons for the WorldView-III and GaoFen2 datasets, as shown in  Figure 3 and  Figure 4. Evaluation on Full-resolution Scene. To assess the performance and generalization capability of our method on full-resolution scenes under real-world conditions, we first trained Pan-LUT on WorldView-II data and then tested it on unseen full-resolution WorldView-II satellite datasets. The real-world dataset consists of 200 newly collected samples from the WorldView-II satellite for evaluation. The results are presented in Table 2. On reduced-resolution scenes, our method falls short of most DNN-based approaches in terms of performance. However, on full-resolution scenes, it outperforms the majority of them when considering the metrics of $D_{\lambda}$, $D_{S}$, and QNR. This demonstrates its strong generalization ability in real-world situations. Additionally, we provide a visual comparison against both traditional and DNN-based methods, as shown in Figure 5. Evaluation on Reduced-resolution Scene. There are two key limitations of DNN-based methods in practical applications. First, real-world remote sensing satellite images typically possess extremely large sizes, thereby incurring substantial computational burdens. Second, computational resources are often limited, which further restricts the applicability of these methods. Therefore, we conduct comparative experiments to evaluate the computational efficiency of various methods on high-resolution images, including $1024\times 1024$, $2048\times 2048$ and $4096\times 4096$ resolutions. For each method, we record the average inference time on 100 images, with all inference steps conducted on a single NVIDIA GeForce RTX 2080Ti GPU for fair comparison. As shown in Table 1, Pan-LUT efficiently processes images at all resolutions. Compared to DNN-based methods, it achieves significantly faster inference speeds, while maintaining comparable speed to traditional methods. Our method easily meets the real-time processing requirements on GPUs, outperforming all other methods by a substantial margin. Notably, only PNN is capable of handling remote sensing satellite images at the $4096\times 4096$ resolution, highlighting the superior efficiency of our approach.
5 Ablation Study
In this section, we conduct several ablation studies on the WorldView-III dataset to investigate the impact of LUT size and evaluate the effectiveness of each proposed component. Size of Look-Up Tables. As shown in Figure 6, changing the LUT size does not lead to a significant drop in performance. This observation suggests that the effectiveness of our proposed method is not dependent on consuming extensive storage resources to increase the LUT size.
First, we explore the effect of PGLUT size, denoted $N_{P}$. Performance improves with larger $N_{P}$, reaching an optimal point at values $N_{P}$ = 9. Beyond this (from 9 to 17), only a small gain of 0.005 dB is observed, while the parameters increase substantially from 236K to 5M, indicating capacity redundancy. We therefore set $N_{P}$ = 9 as the default to balance performance with storage requirements.
For SDLUT, denoted $N_{S}$. we find that increasing $N_{S}$ from 3 to 9 improves performance, but values above 9 cause a slight performance drop.
@@ 5 Ablation Study
Similarly, enlarging AALUT size $N_{A}$ yields only minor gains but substantially increases parameters, especially beyond $N_{A}$ = 9. We set $N_{A}$ = 9 to balance performance with computational efficiency. Effectiveness of Each LUT. We further conduct ablation studies to verify the effectiveness of each LUT. Specifically, we examine the following four settings: (i) only PGLUT, (ii) only SDLUT, (iii) PGLUT and SDLUT, and (iv) SDLUT and AALUT. For (i), the output of PGLUT is directly used as the result. For (ii), the outputs of SDLUT are averaged and then added to the MS image. For (iii), the averaged SDLUT outputs are added to the PGLUT output to generate the result. For (iv), AALUT takes the output of SDLUT as input and performs adaptive aggregation to produce the final result. Results are listed in Table 3. Our observations are as follows: 1) Comparing (i) and (iii), SDLUT effectively captures fine-grained spatial details from the PAN image, thereby enhancing overall performance. 2) Comparing (ii) and (iii), PGLUT provides finer control over spectral transformation, resulting in improved performance. 3) Comparing (ii) and (iv), AALUT demonstrates adaptive aggregation capabilities, effectively learning local contexts and clearly outperforming simple averaging.
Figure 6: Ablation studies on different sizes of the PGLUT, SDLUT and AALUT on the WorldView-III dataset.
Table 3: Ablation study of PGLUT, SDLUT and AALUT on the WorldView-III dataset.

### v2

@@ 3.4 Adaptive Output
where $f_{AOLUT}(\cdot)$ denotes the lookup and interpolation process in the LUT retrieval. More details about AOLUT and the pentalinear interpolation can be found in Section A.4.
@@ 3.5 Loss Function
To achieve satisfying pan-sharpening results, we propose a joint loss for network training. Suppose the batch size is $T$. We first utilize the MSE loss:
$\mathcal{L}_{mse}=\frac{1}{T}\sum_{t=1}^{T}\|HRMS_{t}-GT_{t}\|^{2},$ (7)
@@
$\begin{gathered}\mathcal{L}_{s}=\mathcal{L}_{s}^{PG}+\mathcal{L}_{s}^{SD}+\mathcal{L}_{s}^{AO},\mathcal{L}_{m}=\mathcal{L}_{m}^{PG}+\mathcal{L}_{m}^{SD}+\mathcal{L}_{m}^{AO},\end{gathered}$ (8)
where $\mathcal{L}_{s}^{PG}$, $\mathcal{L}_{s}^{SD}$, and $\mathcal{L}_{s}^{AO}$ denote the smoothness regularizations for PGLUT, SDLUT, and AOLUT, while $\mathcal{L}_{m}^{PG}$, $\mathcal{L}_{m}^{SD}$, and $\mathcal{L}_{m}^{AO}$ represent the monotonicity regularizations for PGLUT, SDLUT, and AOLUT, respectively.
Taking SDLUT as an example, the smoothness regularization can be defined as:
[…]
where $N$ represents the number of bins in each dimension of the LUT. $O_{(i,j,k,m)}$ is the corresponding output for the defined sampling point $(i,j,k,m)$ in LUT. The definitions of $\mathcal{L}{s}^{PG}$ and $\mathcal{L}{s}^{AO}$ are similar to those in Equation 36.
The monotonicity regularization in AOLUT can be defined as:
[…]
where $g(\cdot)$ denotes the ReLU activation function. Similarly, $\mathcal{L}_{m}^{PG}$ and $\mathcal{L}_{m}^{AO}$ are defined in the same way as in Equation 37.
@@
where the two constant parameters $\lambda_{s}$ and $\lambda_{m}$ are used to control the effects of the smoothness and monotonicity regularization terms, respectively. In our experiments, we empirically set $\lambda_{s}$ = 0.0001 and $\lambda_{m}$ = 10. More details about the loss functions can be found in Section A.5
@@ 4.1 Datasets
Remote sensing datasets from three satellites are used in our experiments, including WorldView-II (WV2), GaoFen2 (GF2) and WorldView-III (WV3). Due to the absence of high-resolution multispectral ground truth images in these datasets, we generate the training set using the Wald protocol tool waldtool. Specifically, given the original MS image and its corresponding high-resolution PAN image, they are downsampled by a factor of $r$ to obtain image pairs of MS and PAN, with $r$ set to 4. During training, the original high-resolution MS image is treated as the ground truth, while the MS and PAN images serve as the input image pairs.
4.2 Implementation Details
We compare the proposed Pan-LUT model against several pan-sharpening methods on reduced-resolution scenes from WV2, WV3, and GF2 datasets. Specifically, we choose four traditional pan-sharpening techniques: Brovey brovey, IHS ihs, SFIM liu2000smoothing and GS GS, along with ten deep learning-based approaches: PNN pnn, PanNet pannet, MSDCNN msdcnn, Pan-GAN pangan, SFDI sfiin, UCGAN ucgan, PanFlow yang2023panflownet, PSCINN wang2024panpsci, Pan-Mamba panmamba and TA-DiffHQP tdiff. Several widely used image quality assessment metrics are employed to evaluate the performance of the algorithm, including peak signal-to-noise ratio (PSNR) psnr, structural similarity index (SSIM) ssim, spectral angle mapper (SAM) sam, relative dimensionless global error in synthesis (ERGAS) ergas, spectral distortion index($D_{\lambda}$), spatial distortion index ($D_{S}$) and the quality with no reference (QNR) qnr. The PyTorch framework is implemented in our experiment. During the training phase, we employ an ADAM optimizer with $\beta_{1}=\,\,0.9$ and $\beta_{2}=\,\,0.999$, to update the network parameters for 1000 epochs with a batch size of 1. The learning rate is initialized with $5\times 10^{-4}$. In parallel, a StepLR learning rate adjustment strategy is employed to reduce the learning rate by half after every 200 iterations. The sizes of PGLUT, SDLUT and AOLUT are set to 9, 9 and 9, respectively.
@@ 4.2 Implementation Details
$D_{\lambda}\downarrow$ 0.1026 0.1110 0.1057 0.1063 0.0987 0.1034 0.1042 0.0966 0.0967 0.0966 0.0953  0.0571
$D_{S}\downarrow$ 0.1409 0.1556 0.1446 0.1443 0.1312 0.1305 0.1476 0.1274 0.1271 0.1272 0.1129  0.0640
@@
4.3 Comparison with Other Methods
Evaluation on Reduced-resolution Scene. The quantitative results across three datasets are presented in Table 1, with the best results highlighted in red. Compared to traditional methods, Pan-LUT achieves an average PSNR improvement of 5dB, 7dB, and 7dB across the three datasets, while maintaining inference speeds comparable to those of traditional methods. It is worth noting that the proposed Pan-LUT does not incorporate any network structure, yet it outperforms some DNN-based methods, such as PanNet and PNN, in terms of both performance and inference time. We also provide visual comparisons for the WV3 datasets, as shown in Figure 3.
Evaluation on Full-resolution Scene. To assess the performance and generalization capability of our method on full-resolution scenes under real-world conditions, we first trained Pan-LUT on the reduced-resolution WorldView-II data and then tested it on unseen full-resolution WorldView-II satellite datasets. The real-world dataset consists of 200 newly collected samples from the WorldView-II satellite for evaluation. The results are presented in Table 2. On reduced-resolution scenes, our method falls short of most DNN-based approaches in terms of performance. However, on full-resolution scenes, it outperforms all of them when considering the metrics of $D_{\lambda}$, $D_{S}$, and QNR. This demonstrates its strong generalization ability in real-world situations. Additionally, we provide a visual comparison against both traditional and DNN-based methods, as shown in Figure 4.
Computation Efficiency Comparison. We conduct three experiments to comprehensively evaluate the computational efficiency of all methods: (1) testing the maximum image size that each method can handle on 11GB and 24GB GPUs; (2) measuring their inference time on the CPU; and (3) evaluating their inference time on an RTX 2080 Ti GPU with $\times2K$ and $\times4K$ images. For each method, we record the average inference time on 100 images. As shown in Figure 1, even with a 24GB GPU, existing methods fail to process $\times8K$ images, while our method can handle $\times9K$ images on an 11GB GPU. In environments without GPU acceleration, most methods exhibit unsatisfactory inference speed. As shown in Tabel 1, Pan-LUT efficiently processes images at all resolutions. Compared to DNN-based methods, it achieves significantly faster inference speeds, while maintaining comparable speed to traditional methods. Our method easily meets the real-time processing requirements on GPUs, outperforming all other methods by a substantial margin. Notably, only PNN is capable of handling remote sensing satellite images at the $\times4K$ resolution, highlighting the superior efficiency of our approach.
4.4 Ablation Study
Figure 5: Ablation studies on different sizes of the PGLUT, SDLUT and AOLUT on the WorldView-III dataset.
Table 3: Memory requirements (K).
Module 3 6 9 17 33
AOLUT 0.97 31.10 236.20 5679.43 156541.57
SDLUT 0.08 1.30 6.56 83.52 1185.92
PGLUT 1.21 38.88 295.24 7099.28 195676.96
Table 4: Ablation study of PGLUT, SDLUT and AOLUT on the WorldView-III dataset.
@@ 4.4 Ablation Study
(i) only PGLUT 25.2788 0.7742 0.1130 5.4852
(ii) only SDLUT 26.1433 0.8165 0.1123 5.4211
(iii) only AOLUT 28.9554 0.8677 0.1001 3.7661
(iv) PGLUT and SDLUT 29.1315 0.8795 0.0988 3.6785
(v) SDLUT and AOLUT 29.2754 0.9010 0.0934 3.6033
(vi) PGLUT and AOLUT 29.4875 0.8925 0.0900 3.4977
$\star$ Pan-LUT (Ours) 29.7376 0.9106 0.0815 3.3934
Size of Look-Up Tables. As shown in Figure 5, changing the LUT size does not lead to a significant drop in performance. This observation suggests that the effectiveness of our proposed method is not dependent on consuming extensive storage resources to increase the LUT size. First, we examine the effect of PGLUT size, denoted $N_{P}$. Performance improves with larger $N_{P}$, reaching an optimal point at values $N_{P}$ = 9. Beyond this (from 9 to 17), only a minor gain of 0.01 dB is observed, while the number of parameters increases substantially from 236K to 5M, indicating capacity redundancy. Therefore, we set $N_{P}$ = 9 as the default to balance performance with storage requirements. For SDLUT, denoted $N_{S}$, increasing $N_{S}$ from 3 to 9 improves performance, but values above 9 cause a slight performance drop. Similarly, enlarging the AOLUT size $N_{A}$ yields only minor gains but substantially increases parameters, especially beyond $N_{A}$ = 9. We thus set $N_{A}$ = 9 to balance performance with computational efficiency. We provide the parameter count for each LUT of different sizes in Tabel 5, which can be calculated as follows:
$\begin{gathered}Param_{PGLUT}=5N^{5},Param_{SDLUT}=N^{4},Param_{AOLUT}=4N^{5}.\end{gathered}$ (12)
@@
Effectiveness of Each LUT. We further conduct ablation studies to verify the effectiveness of each LUT. Results are listed in Table 5. Our observations are as follows: 1) Comparing (i) with (iv) and (iii) with (v), SDLUT effectively captures fine-grained spatial details from the PAN image, thereby enhancing overall performance. 2) Comparing (ii) with (iv) and (iii) with (iv), PGLUT provides finer control over spectral transformation, resulting in improved performance. 3) Comparing (ii) with (v) and (i) with (vi), AOLUT demonstrates adaptive aggregation capabilities.
5 Conclusion
In this paper, we propose a novel learnable LUT framework, called Pan-LUT, which strikes an optimal balance between performance and computational efficiency for high-resolution remote sensing images in pan-sharpening. The proposed method makes it possible to process $15K\times 15K$ remote sensing images on a 24GB GPU and processes a $\times9K$ image in under 1 ms using one RTX 2080 Ti GPU. Extensive experiments on various satellite datasets demonstrate the effectiveness and efficiency of Pan-LUT.
Appendix A Appendix
The following contents are provided in the appendix:
  • Limitation and Border Impact.
  • More Technical Details.
  • More ablation study.
  • More comparison experiments.
  • More visualization about our experiments.
A.1 Limitation and Border Impact
Limitation. One limitation of existing methods is the lack of neural network integration. While combining LUTs with neural structures may offer greater modeling capacity, it often comes at the cost of computational efficiency. Conversely, LUTs alone require substantial storage space, making them less suitable for high-dimensional input processing. The space complexity of an n-dimensional LUT is $O(ND^{n})$, meaning its size increases exponentially with higher dimensions. This exponential growth severely limits its practical applicability, particularly when dealing with high-dimensional inputs. Lightweight network architectures must be further explored to facilitate efficient integration with LUTs. Meanwhile, more efficient LUT designs are essential for processing high-dimensional inputs.
A.2 Details of PGLUT
Input: Suppose that the MS image ($MS\in R^{H/r\times W/r\times C}$) and the PAN image ($P\in R^{H\times W\times 1}$) are given, the input of PGLUT ($I_{PG}\in R^{H\times W\times(C+1)}$) can be formulated as:
$I_{PG}=Concat(P,MS\uparrow),$ (13)
where $Concat(\cdot)$ is the concatenation operation. $MS\uparrow$ denotes the upsampled MS image. $MS\uparrow$ and $P$ have the same spatial resolution.
Output: the output of PGLUT ($O_{PG}\in R^{H\times W\times(C+1)}$) can be formulated as:
$O_{PG}=F_{PGLUT}(I_{PG}),$ (14)
where $F_{PGLUT}(\cdot)$ denotes the lookup and pentalinear interpolation based on the PGLUT.
PAN-guided indexing strategy. In a MS image, pixels from different spatial locations may have identical values (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j}$, where $i\neq j$). The LUT maps these identical inputs to the same output. The PAN-guided indexing strategy provides a more flexible indexing mechanism for the LUT. Specifically, it additionally considers the pixels at corresponding spatial positions in the PAN image (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j},pa_{i}\neq pa_{j}$, where $i\neq j$), thereby achieving finer-grained mapping. Given an input value $I_{(w,h)}=\{pa_{(w,h)}^{I},r_{(w,h)}^{I},g_{(w,h)}^{I},b_{(w,h)}^{I},nir_{(w,h)}^{I}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image, the input index to the PGLUT based on the input value can be represented as (x, y, z, s, e). PGLUT first performs a lookup operation to locate the nearest 32 adjacent elements around the input index in the PGLUT. We use $(i,j,k,m,n)$ to denote the coordinates of a defined sampling point in PGLUT, which can be calculated as follows:
$\begin{gathered}x=\frac{pa_{(w,h)}^{I}}{V_{max}}\cdot N,y=\frac{r_{(w,h)}^{I}}{V_{max}}\cdot N,z=\frac{g_{(w,h)}^{I}}{V_{max}}\cdot N,\\ s=\frac{b_{(w,h)}^{I}}{V_{max}}\cdot N,e=\frac{nir_{(w,h)}^{I}}{V_{max}}\cdot N,\\ i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,n=\lfloor e\rfloor,\end{gathered}$ (15)
where $V_{max}$ denotes the maximum value (e.g., 255, 1023 or 2047). $\lfloor\cdot\rfloor$ denotes the floor function. Then, the offset between the input precise index $(x,y,z,s,e)$ and the computed sampling point $(i,j,k,m,n)$ can be computed:
$\begin{gathered}d_{x}=x-i,d_{y}=y-j,d_{z}=z-k,\\ d_{s}=s-m,d_{e}=e-m,\\ d_{-x}=1-d_{x},d_{-y}=1-d_{y},d_{-z}=1-d_{z},\\ d_{-s}=1-d_{s},d_{-e}=1-d_{e}.\end{gathered}$ (16)
Pentalinear Interpolation. After locating 32 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$\begin{gathered}O_{PG(x,y,z,s,e)}=d_{-x}d_{-y}d_{-z}d_{-s}d_{-e}O_{(i,j,k,m,n)}+d_{x}d_{-y}d_{-z}d_{-s}d_{-e}O_{(i+1,j,k,m,n)}\\ +d_{-x}d_{y}d_{-z}d_{-s}d_{-e}O_{(i,j+1,k,m,n)}+d_{-x}d_{-y}d_{z}d_{-s}d_{-e}O_{(i,j,k+1,m,n)}\\ +d_{-x}d_{-y}d_{-z}d_{s}d_{-e}O_{(i,j,k,m+1,n)}+d_{-x}d_{-y}d_{-z}d_{-s}d_{e}O_{(i,j,k,m,n+1)}\\ +d_{x}d_{y}d_{-z}d_{-s}d_{-e}O_{(i+1,j+1,k,m,n)}+d_{x}d_{-y}d_{z}d_{-s}d_{-e}O_{(i+1,j,k+1,m,n)}\\ +d_{x}d_{-y}d_{-z}d_{s}d_{-e}O_{(i+1,j,k,m+1,n)}+d_{x}d_{-y}d_{-z}d_{-s}d_{e}O_{(i+1,j,k,m,n+1)}\\ +d_{-x}d_{y}d_{z}d_{-s}d_{-e}O_{(i,j+1,k+1,m,n)}+d_{-x}d_{y}d_{-z}d_{s}d_{-e}O_{(i,j+1,k,m+1,n)}\\ +d_{-x}d_{y}d_{-z}d_{-s}d_{e}O_{(i,j+1,k,m,n+1)}+d_{-x}d_{-y}d_{z}d_{s}d_{-e}O_{(i,j,k+1,m+1,n)}\\ +d_{-x}d_{-y}d_{z}d_{-s}d_{e}O_{(i,j,k+1,m,n+1)}+d_{-x}d_{-y}d_{-z}d_{s}d_{e}O_{(i,j,k,m+1,n+1)}\\ +d_{x}d_{y}d_{z}d_{-s}d_{-e}O_{(i+1,j+1,k+1,m,n)}+d_{x}d_{y}d_{-z}d_{s}d_{-e}O_{(i+1,j+1,k,m+1,n)}\\ +d_{x}d_{y}d_{-z}d_{-s}d_{e}O_{(i+1,j+1,k,m,n+1)}+d_{x}d_{-y}d_{z}d_{s}d_{-e}O_{(i+1,j,k+1,m+1,n)}\\ +d_{x}d_{-y}d_{z}d_{-s}d_{e}O_{(i+1,j,k+1,m,n+1)}+d_{x}d_{-y}d_{-z}d_{s}d_{e}O_{(i+1,j,k,m+1,n+1)}\\ +d_{-x}d_{y}d_{z}d_{s}d_{-e}O_{(i,j+1,k+1,m+1,n)}+d_{-x}d_{y}d_{z}d_{-s}d_{e}O_{(i,j+1,k+1,m,n+1)}\\ +d_{-x}d_{y}d_{-z}d_{s}d_{e}O_{(i,j+1,k,m+1,n+1)}+d_{-x}d_{-y}d_{z}d_{s}d_{e}O_{(i,j,k+1,m+1,n+1)}\\ +d_{x}d_{y}d_{z}d_{s}d_{-e}O_{(i+1,j+1,k+1,m+1,n)}+d_{x}d_{y}d_{z}d_{-s}d_{e}O_{(i+1,j+1,k+1,m,n+1)}\\ +d_{x}d_{y}d_{-z}d_{s}d_{e}O_{(i+1,j+1,k,m+1,n+1)}+d_{x}d_{-y}d_{z}d_{s}d_{e}O_{(i+1,j,k+1,m+1,n+1)}\\ +d_{-x}d_{y}d_{z}d_{s}d_{e}O_{(i,j+1,k+1,m+1,n+1)}+d_{x}d_{y}d_{z}d_{s}d_{e}O_{(i+1,j+1,k+1,m+1,n+1)},\end{gathered}$ (17)
where $O_{(i,j,k,m,n)}$ represents the value of the LUT at the coordinate $(i,j,k,m,n)$.
A.3 Details of SDLUT
Input: The input image ($O_{PG}\in R^{H\times W\times{(C+1)}}$) of SDLUT is the output of PGLUT.
Output: The output of SDLUT ($O_{SD}\in R^{H\times W\times{(C+1)}}$) can be formulated as:
${O_{SD}}=F_{SDLUT}(O_{PG}),\\$ (18)
where $F_{SDLUT}(\cdot)$ denotes the lookup and quadrilinear interpolation based on the SDLUT.
Look Up. Specifically, for each channel of input, SDLUT operates by iterating through pixels one at a time, treating each as the current pixel during processing. Given the input current pixel and ajacent pixels, which denoted as $p_{(w,h)}$, $p_{(w+1,h)}$, $p_{(w,h+1)}$ and $p_{(w+1,h+1)}$ respectively, the input index to the SDLUT based on the input can be represented as $(x,y,z,s)$. SDLUT first performs a lookup operation to locate the nearest 16 adjacent elements around the input index in the SDLUT. We use $(i,j,k,m)$ to denote the coordinates of a defined sampling point in PGLUT, which can be calculated as follows:
$\begin{gathered}x=\frac{p_{(w,h)}}{V_{max}}\cdot N,y=\frac{p_{(w+1,h)}}{V_{max}}\cdot N,\\ z=\frac{p_{(w,h+1)}}{V_{max}}\cdot N,s=\frac{p_{(w+1,h+1)}}{V_{max}}\cdot N,\\ i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,\end{gathered}$ (19)
where $V_{max}$ denotes the maximum pixel value. $\lfloor\cdot\rfloor$ signifies the floor function. Then, the offset between the input precise index $(x,y,z,s)$ and the computed sampling point $(i,j,k,m)$ can be computed:
$\begin{gathered}d_{x}=x-i,d_{y}=y-j,d_{z}=z-k,d_{s}=s-m,\\ d_{-x}=1-d_{x},d_{-y}=1-d_{y},\\ d_{-z}=1-d_{z},d_{-s}=1-d_{s}.\end{gathered}$ (20)
Quadrilinear Interpolation. After locating 16 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$\begin{gathered}O_{SD(x,y,z,s)}=d_{-x}d_{-y}d_{-z}d_{-s}O_{(i,j,k,m)}+d_{x}d_{-y}d_{-z}d_{-s}O_{(i+1,j,k,m)}\\ +d_{-x}d_{y}d_{-z}d_{-s}O_{(i,j+1,k,m)}+d_{-x}d_{-y}d_{z}d_{-s}O_{(i,j,k+1,m)}\\ +d_{-x}d_{-y}d_{-z}d_{s}O_{(i,j,k,m+1)}+d_{x}d_{y}d_{-z}d_{-s}O_{(i+1,j+1,k,m)}\\ +d_{x}d_{-y}d_{z}d_{-s}O_{(i+1,j,k+1,m)}+d_{x}d_{-y}d_{-z}d_{s}O_{(i+1,j,k,m+1)}\\ +d_{-x}d_{y}d_{z}d_{-s}O_{(i,j+1,k+1,m)}+d_{-x}d_{y}d_{-z}d_{s}O_{(i,j+1,k,m+1)}\\ +d_{-x}d_{-y}d_{z}d_{s}O_{(i,j,k+1,m+1)}+d_{x}d_{y}d_{z}d_{-s}O_{(i+1,j+1,k+1,m)}\\ +d_{x}d_{y}d_{-z}d_{s}O_{(i+1,j+1,k,m+1)}+d_{x}d_{-y}d_{z}d_{s}O_{(i+1,j,k+1,m+1)}\\ +d_{-x}d_{y}d_{z}d_{s}O_{(i,j+1,k+1,m+1)}+d_{x}d_{y}d_{z}d_{s}O_{(i+1,j+1,k+1,m+1)},\\ \end{gathered}$ (21)
where $O_{(i,j,k,m)}$ represents the value of the LUT at the coordinate $(i,j,k,m)$.
Rotation-indexing strategy. In general, the performance of the SDLUT can be improved when more pixels are considered. In order to exploit more area in the image, we use a Rotation-indexing strategy in the training phase. For our SDLUT, 4 rotational ensemble with 0, 90, 180, and 270 degrees covers total $3\times 3$ pixels. Each output from the 4 rotations is defined as follow: